OpenStudy (grace111):

IMPORTANT NOTE : WILL GIVE MEDAL AND FAN FOR BEST ANSWER !!! x QUESTION 1 ) Circle TWO different numbers which multiply together to make 1 million . 50 200 5000 20 000 100 000 QUESTION 2 ) Circle ALL the multiples of 9 in the list of numbers. 19 29 36 51 81. Thanks guys xxx

OpenStudy (grace111):

:( can someone plzz try to help me solve this? I will really reward the very best a medal and fan

OpenStudy (grace111):

Its 500000

OpenStudy (anonymous):

Yep :)

OpenStudy (grace111):

@ღChloeღ is this the answer ?

OpenStudy (anonymous):

9, 18, 27, 36, 45, 54, 45, 63, 72, 81, 90

OpenStudy (anonymous):

No its not that would be 1 billion

undeadknight26 (undeadknight26):

oh oops.

OpenStudy (anonymous):

We all make mistakes

OpenStudy (grace111):

so what is the answer for question 1 ?

OpenStudy (anonymous):

So far I see 2 multiples of 9. Who about you?

OpenStudy (grace111):

and question 2?

OpenStudy (anonymous):

9, 18, 27, 36, 45, 54, 45, 63, 72, 81, 90<---multiples of 9

OpenStudy (grace111):

oh thank you @ღChloeღ :) hmm.. but Im not sure about question 1 it says that you have to circle TWO DIFFERENT numbers which multiply together to make 1 million. could you explain a bit more only question 1

OpenStudy (grace111):

but anyway you will get the medal and fan :) ! x

OpenStudy (anonymous):

Thanks :)

OpenStudy (grace111):

oh thank you very much for your help ! your a good helper

OpenStudy (grace111):

:)

OpenStudy (anonymous):

^_^

OpenStudy (radar):

20,000 = 2 10^4 50 = 5 10^1 (20,000)(50) = 2 10^4 X 5 10^1 = 10 10^5 = 10^6 = 1 million.

OpenStudy (radar):

Look it over and decide for yourself. and how does 20,000 X 5,000 equals 1,000,000? 2 10^4 X 5 10^3 = 10 10^7 = 10^8 or MORE than 1 million. Think.

OpenStudy (radar):

I chose from the list of numbers that were listed. The first one and the fourth one @eChloe

OpenStudy (radar):

What list were you looking at. Please in the future, show the method you use and the basis not just an answer, it could be helpful to @grace111

OpenStudy (radar):

Yes, you did help, please disregard that comment. But I think you gave her an incorrect answer.

OpenStudy (radar):

Or at least an incorrect choice.

OpenStudy (radar):